A 3 kg ball rests on top of a spring on the ground and the weight of the ball compresses the spring a distance of 3 cm. What is the spring constant of the spring?


1
Expert's answer
2021-01-10T18:27:03-0500

We can find the spring constant of the spring from the Hooke's law:


"F=kx,""mg=kx,""k=\dfrac{mg}{x}=\dfrac{3\ kg\cdot 9.8\ \dfrac{m}{s^2}}{0.03\ m}=980\ \dfrac{N}{m}."

Answer:

"k=980\ \dfrac{N}{m}."